The ratio of men to women working for a company is 3 to 2 . If there are 32 women working for the company, what is the total number of employees?
Answer by harpazo(655)

3/2 = x/32
3 times 32 = 2x
96 = 2x
96/2 = x
48 = x
The total number of employees is 32 women plus 48 men or 80.
Done!